Suppose a celebrity gains a twitter follower every 15 minutes, with a pattern that conforms to the exponential distribution. He has 999,999 followers. What is the probability he'll have to wait at least half an hour to reach 1 million?
Answer #1
here for exponential distribution parameter β = 15

probability he'll have to wait at least half an hour to reach 1 million =P(no new follower in next 30 minutes):

P(X>30)=1-P(X<30)=1-(1-exp(-30/15))= 0.1353